the sum of two numbers is 62. the smaller number is 22 less than the larger number. what are the two numbers?
Answer by stanbon(57214)

the sum of two numbers is 62. the smaller number is 22 less than the larger number. what are the two numbers?
----------
Equations:
x + y = 62
x = y - 22
-----
Substitute for "x" and solve for "y":
y-22 + y = 62
2y = 84
y = 42
----
Solve for "x":
x = y - 22
x = 42-22
x = 20
============
Cheers,
Stan H.
============